9. For the following set of scores, find the value of each expresion: a.Σ b. (EX)2 c. Σ(X-1) d. Σ(x-1)?
Answer #1

Q)

X X2 (X - 1) (X - 1)2
3 9 (3 - 1) = 2 22 = 4
2 4 (2 - 1) = 1 12 = 1
5 25 (5 - 1) = 4 42 = 16
1 1 (1 - 1) = 0 0
3 9 (3 - 1) = 2 22 = 4
X = 14 X2-48 sum (X-1) = 9 (X- 1) 25

a) X2-48

b) (〉 X)2= (14)2-196

ext { c) } sum (X -1) = 9

d) 〉 (X-1)2-25